OpenStudy (anonymous):

in standard . m x 5 + 4(m + 5/9

OpenStudy (sidsiddhartha):

standard form is-- \[y=a(x-h)^2+k\]

OpenStudy (sidsiddhartha):

so just take the -11 to other side

OpenStudy (anonymous):

how?

OpenStudy (sidsiddhartha):

\[y-11=-2(x-2)^2\\y=-2(x-2)^2+11\] thats the standard form

OpenStudy (sidsiddhartha):

ok?

OpenStudy (sidsiddhartha):

yes now simplify it

OpenStudy (sidsiddhartha):

\[y=-2(x-2)^2+11=-2(x^2-4x+4)+11=-2x^2+8x+3\]

OpenStudy (anonymous):

ohhh ok

OpenStudy (sidsiddhartha):

yup :)